Replace each letter with a number to make ths problem correct. Identical letters shoul be replaced with identical numbers.
P A R T S
* 4
__________
= S T R A P
hint: Since 4 * p =a 1-digit answer, p must equal 1 or 2. Remember 4* any other letter can give either a 1- or 2-digit answer.

Answer by Alan3354(69443)

P A R T S
* 4
__________
= S T R A P
hint: Since 4 * p =a 1-digit answer, p must equal 1 or 2. Remember 4* any other letter can give either a 1- or 2-digit answer.
-------------------
And P has to be an even number, since 4*S = P. So P = 2
2 A R T S
* 4
__________
= S T R A 2
S has to be either 3 or 8, since 4*S = -2. 3 is out, since the LS digit of the product is S, so S = 8.
2 A R T 8
* 4
__________
= 8 T R A 2
Since nothing was carried from the 4*A, and 2 is used, A has to be 1.
2 1 R T 8
* 4
__________
= 8 T R 1 2
For the product to end in 12, T 8 has to be either 28 or 78. 2 is used, so T = 7
2 1 R 7 8
* 4
__________
= 8 7 R 1 2
R*4 has to be 30 something, since 4*21 is 84, so a 3 had to be carried. Only 8 or 9 fills that, and 8 is taken. R = 9
2 1 9 7 8
* 4
__________
= 8 7 9 1 2